OpenStudy (anonymous):

I DO NOT JUST WANT THE ANSWER i want to learn can anyone give me a formula for this question and a brief steps. HELP PLEASE! A bag contains 6 cherry, 3 orange, and 2 lemon candies. You reach in and take 3 pieces of candy at random. The probability that you have 1 cherry candy and 2 lemon candies

OpenStudy (anonymous):

Well, that's good. You don't have to put it in all-caps, though. So, here: The probability of an event \(E\) is: \[ P(E)=\frac{\text{desired outcomes}}{\text{possible outcomes}} \]The probability of multiple, independent events \(A,B,...\) all happening (that is, the outcome of one event does not change the probability of the other) is: \[ P(A\cap B\cap\cdots)=P(A)\cdot P(B)\cdots \]So, we are going to have that (1) the probability of picking a cherry for your first time is: \[ P(A)=\frac{6}{6+3+2}=\frac{6}{11} \]The next time you pick, you now are missing one candy from the bag, so we have, for one lemon: \[ P(B)=\frac{2}{5+3+2}=\frac{1}{5} \]Now, we're missing one lemon and a cherry, so our final pick will have a probability of: \[ P(C)=\frac{1}{5+3+1}=\frac{1}{9} \]Now, we wish to find the probability of ALL of these events, that is: \[ P(A\cap B\cap C)=P(A)\cdot P(B)\cdot P(C) \]What would this be?

OpenStudy (anonymous):

6/11*2/10*1/9=12/990=0.01212=1.2%?

OpenStudy (anonymous):

That is correct.

OpenStudy (anonymous):

thank you!